Fermi National Laboratory


Accelerator Update for April 6, 2005 - April 8, 2005

Wednesday April 6
The day shift began with Operations monitoring store 4078 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE and the Meson MCenter experiments taking beam.

At 9:36 AM, a Booster quadrupole (B:Q802) power supply tripped off for the first of four times in thirty-one minutes. Operations notified experts.

Operations put the Recycler into shot setup at 12:43 PM to stash antiprotons.

At 1:25 PM, B:Q802 tripped off. The Booster expert found a metering problem with this power supply and also the supply for DOG3. He repaired both and stacking resumed at 1:58 PM.

MCenter resumed taking beam at 2:01 PM.

Tevatron (TeV) experts flew wires at 4:12 PM.

Operations established beam to the Meson MTest experiment at 4:27 PM.

At 6:04 PM, Pbar suffered a loss of antiproton cooling. A number of Debuncher amplifiers tripped off and wouldn't reset. Operators turned off the Accumulator at 6:18 PM due to Debuncher cooling problems messing up the beam. They called in an expert to investigate.

Operators put the Recycler into shot setup at 7:08 PM.

A Debuncher expert found a power supply with a blown fuse at 8:11 PM.

Operations began antiproton transfers at 8:18 PM. There were six transfers, but due to problems Recycler only got 36mA.

Pbar resumed stacking at 8:53 PM.

At 11:06 PM, the TeV quenched due to a separator spark at sector A49. Operations notified experts.

Thursday April 7
The midnight shift began with TeV recovering from a quench, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations put the TeV through a dry squeeze at 1:05 AM.

Operations put the TeV into shot setup at 2:23 AM.

At 3 AM, Operations reported that shot setup was on hold while operators investigated more separator problems. They replaced a control card.

Operations put the TeV into a dry squeeze at 4:16 AM.

Operations put the TeV into shot setup at 4:38 AM.

Operations established combination store 4081 at 6:32 AM with an initial luminosity of 101.01E30. Pbar resumed stacking.

MCenter resumed taking beam at 6:53 AM.

MCenter experiments reported at 10:10 AM that they needed their beam disabled for scheduled work.

Operations established beam to MTest at 11:23 AM.

Operations put the Recycler into shot setup at 12:29 PM.

The Recycler transfers ended and Pbar stacking resumed at 1:44 PM.

MTest resumed taking beam at 1:55 PM.

At 2:07 PM, a Booster power supply (GMPS) tripped off on a ground fault indication. Operators began preparing Booster for an access 2:20 PM.

Operators reported finding a Booster LCW hose leak on a candy cane. Water technicians were called to make repairs. The techs found two more leaks. leak

Water techs reported at 5:29 PM that the Booster LCW leaks were fixed.

Pbar resumed stacking at 5:39 PM.

MiniBooNE, MCenter, and MTest all resumed taking beam at 5:41 PM.

Operations put the Recycler into shot setup at 11:01 PM.

Friday April 8
The midnight shift began with Operations monitoring store 4081 and the stash, with Booster having beam problems, and with the Recycler shot delayed.

Operations reported at 1:17 AM that they found a Booster corrector power supply not operating properly.

MCenter resumed stacking at 1:28 AM.

At 1:55 AM, a Booster RF station (BRF16) tripped off for the first of five times in six hours.

Operations began stashing antiprotons into the Recycler at 6:02 AM.

The Recycler took seven transfers and then Pbar resumed stacking at 6:52 AM.

MCenter resumed taking beam at 7:09 AM.

The Plan for Friday
The plan for today is to terminate store 4081 around 10 AM. The Run Coordinator will allow the experiments, machines, and cryo system techs to conduct two hours of work. The next TeV shot setup, which should begin a little after noon, will be a Recycler only store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/8/2005   email Fermilab